# Auditing Student–AI Collaboration: A Case Study of Online Graduate CS Students

### v2

@@ 6. Discussion
Across RQ1–RQ3, our findings collectively reveal that students’ engagement with AI in academic contexts is neither uniformly enthusiastic nor uniformly cautious. Instead, students demonstrate a nuanced, task-sensitive reasoning process that balances desired automation, actual usage, perceived benefits, and anticipated risks. In this section, we synthesize findings from all three research questions to discuss (1) how gaps between desired and actual automation reflect students’ conditional trust in AI systems (RQ1), (2) how motivations and concerns are selectively activated based on task characteristics (RQ2), and (3) how students’ articulated design expectations respond directly to these tensions (RQ3). Together, these insights highlight important implications for the design of educational AI systems that aim to support learning without undermining student agency.
@@ 6.1. Conditional Automation and the Limits of Adoption
From a broader perspective, RQ1 indicates that students are actively negotiating the boundary between assistance and delegation. Rather than seeking maximal automation, students appear to prefer retaining human control in tasks where outcomes carry social, intellectual, or evaluative significance. This finding challenges narratives that frame AI adoption as a linear progression toward increasing automation, and instead positions student–AI interaction as a calibrated and situational process.
@@ 6.2. Task-Specific Motivations and Risk Awareness
Taken together, RQ2 shows that students are not merely efficiency seekers. They are sensitive to the cognitive role of the task and actively evaluate whether AI use supports or threatens their learning objectives. This sensitivity provides important context for understanding the adoption gaps observed in RQ1.
@@ 6.3. Design Expectations as Responses to Identified Risks
Explainability and opportunities for user intervention further reinforce this orientation. Students expressed interest in AI systems that expose reasoning steps and allow mid-course correction, not to eliminate effort, but to support active engagement and learning. Notably, even explicitly skeptical participants did not reject AI outright; instead, they emphasized the importance of designs that respect user agency and avoid creating an illusion of infallibility.
6.4. Implications for Educational AI Design
Synthesizing across all three research questions, our findings suggest that effective educational AI systems should prioritize calibrated assistance rather than maximal automation. Students’ desired future use of AI is not limited by lack of interest, but by concerns about trust, learning, and accountability.
Designs that foreground transparency, uncertainty communication, and explainability may help close the gap between desired and actual AI usage observed in RQ1, particularly for tasks where students currently hesitate. More broadly, our results indicate that supporting learning requires AI systems that scaffold thinking, invite verification, and preserve students’ role as primary decision-makers. Educational AI that emphasizes efficiency alone risks undermining precisely those cognitive processes students seek to protect.
7. Limitations
This study has several limitations that should be considered when interpreting the results.
First, the sample size of this study is relatively small and drawn from a single disciplinary population, namely computer science (CS) students. While CS students are typically early adopters of AI tools and are more familiar with interacting with LLM-based systems, their experiences and expectations may not generalize to students from other academic disciplines. Students in the humanities, social sciences, or professional fields may engage with AI differently, particularly in terms of task types, norms of academic integrity, and tolerance for automation. As a result, the patterns observed in desired automation, usage levels, and concerns may reflect discipline-specific practices rather than universal student perspectives. Future work should extend this investigation to a more diverse student population to examine how disciplinary context shapes student–AI interaction.
Second, unlike prior work (Shao et al., 2025), which incorporates expert annotation to assess the objective capabilities of AI systems on specific tasks, our study relies on self-reported measures of AI usage and perceived automation levels. This means that our analysis reflects students’ subjective experiences and judgments rather than externally validated performance metrics. While this approach is well suited for understanding students’ perceptions, trust calibration, and decision-making, it does not allow us to directly compare perceived AI capability with actual system performance. Consequently, mismatches between desired automation and usage should be interpreted as reflecting students’ beliefs and caution, rather than definitive assessments of AI effectiveness. Future research could combine user-reported data with expert evaluation or benchmark-based performance assessments to more fully characterize the relationship between perceived and actual AI capabilities in educational contexts.
Despite these limitations, our study provides valuable insights into how students reason about AI use across academic tasks, highlighting the nuanced trade-offs they make between efficiency, learning, and risk.

### v3

@@ 7. Conclusion
Appendix A Full Participant List
Table 4. Participant-level demographics of OMSCS students who reported AI tool use in coursework (N=57).
ID Age Group OMSCS Specialization AI Usage Frequency
P1 30–40 Human–Computer Interaction Several times per week
P2 25–30 Machine Learning Daily or almost daily
P3 25–30 Human–Computer Interaction Often
P4 18–24 Artificial Intelligence Often
P5 30–40 Human–Computer Interaction Several times per week
P6 25–30 Human–Computer Interaction Rarely
P7 30–40 Artificial Intelligence Several times per week
P8 Above 40 Computing Systems Several times per week
P9 30–40 Artificial Intelligence Daily or almost daily
P10 18–24 Artificial Intelligence Several times per week
P11 18–24 Not specified Daily or almost daily
P12 18–24 Artificial Intelligence Daily or almost daily
P13 18–24 Undecided Often
P14 30–40 Undecided Daily or almost daily
P15 30–40 Human–Computer Interaction Daily or almost daily
P16 30–40 Computing Systems Rarely
P17 18–24 Undecided Several times per week
P18 30–40 Computing Systems Often
P19 25–30 Artificial Intelligence Often
P20 30–40 Human–Computer Interaction Several times per week
P21 25–30 Computing Systems Several times per week
P22 18–24 Human–Computer Interaction Rarely
P23 18–24 Human–Computer Interaction Several times per week
P24 30–40 Artificial Intelligence Daily or almost daily
P25 18–24 Artificial Intelligence Often
P26 25–30 Human–Computer Interaction Rarely
P27 18–24 Human–Computer Interaction Rarely
P28 18–24 Human–Computer Interaction Rarely
P29 30–40 Machine Learning Daily or almost daily
P30 25–30 Human–Computer Interaction Rarely
P31 25–30 Human–Computer Interaction Daily or almost daily
P32 25–30 Computing Systems Daily or almost daily
P33 30–40 Artificial Intelligence Daily or almost daily
P34 18–24 Machine Learning Daily or almost daily
P35 25–30 Human–Computer Interaction Never
P36 30–40 Artificial Intelligence Never
P37 18–24 Artificial Intelligence / Machine Learning Several times per week
P38 Above 40 Artificial Intelligence Daily or almost daily
P39 18–24 Undecided Several times per week
P40 18–24 Human–Computer Interaction Rarely
P41 25–30 Artificial Intelligence Daily or almost daily
P42 18–24 Artificial Intelligence Daily or almost daily
P43 18–24 Not specified Daily or almost daily
P44 25–30 Human–Computer Interaction Rarely
P45 25–30 Human–Computer Interaction Several times per week
P46 18–24 Human–Computer Interaction Rarely
P47 18–24 Artificial Intelligence Often
P48 18–24 Artificial Intelligence Often
P49 30–40 Computing System Daily or almost daily
P50 25–30 Human–Computer Interaction Daily or almost daily
P51 25–30 Human–Computer Interaction Several times per week
P52 Above 40 Machine Learning Daily or almost daily
P53 Above 40 Human–Computer Interaction Rarely
P54 18–24 Artificial Intelligence Several times per week
P55 25–30 Human–Computer Interaction Several times per week
P56 25–30 Human–Computer Interaction Daily or almost daily
P57 30–40 Artificial Intelligence Daily or almost daily